# RankMamba: Benchmarking Mamba’s Document Ranking Performance in the Era of Transformers

### v2

@@ Abstract
Transformer structure has achieved great success in multiple applied machine learning communities, such as natural language processing (NLP), computer vision (CV) and information retrieval (IR). Transformer architecture’s core mechanism—attention requires $O(n^{2})$ time complexity in training and $O(n)$ time complexity in inference. Many works have been proposed to improve the attention mechanism’s scalability, such as Flash Attention and Multi-query Attention. A different line of work aims to design new mechanisms to replace attention. Recently, a notable model structure—Mamba  which is based on state space models, has achieved transformer-equivalent performance in multiple sequence modeling tasks.
In this work, we examine Mamba’s efficacy through the lens of a classical IR task—document ranking. A reranker model takes a query and a document as input, and predicts a scalar relevance score. This task demands the language model’s ability to comprehend lengthy contextual inputs and to capture the interaction between query and document tokens. We find that (1) Mamba models achieve competitive performance compared to transformer-based models with the same training recipe; (2) but also have a lower training throughput in comparison to efficient transformer implementations such as flash attention. We hope this study can serve as a starting point to explore Mamba  models in other classical IR tasks. Our code implementation and trained checkpoints are made public to facilitate reproducibility. [^1] https://github.com/zhichaoxu-shufe/RankMamba
@@ 1 Introduction
Transformer structure has been the status quo within NLP and IR community since the introduction of Transformers [Vas+17] and BERT [Dev+19]. Compared to the classical RNN structure, the transformer structure is better at capturing long-range dependencies and also can be pre-trained on a large scale. However, in inference time, the transformer structure requires $O(n)$ time complexity and $O(n\cdot d)$ space complexity to store KV-cache, where $n$ denotes sequence length and $d$ denotes the size of the hidden state, making it less efficient compared to RNN. In addition, the transformer structure is also hard to extrapolate to context length longer than its pre-training length.
Recently, there has been a growing interest in developing alternative structures for modeling sequence data. For example, RWKV [Pen+23] combines the efficient parallelizable training of transformers with the efficient inference of RNNs. Another notable structure is state space models (SSM) [GD23], which are related to CNN, RNN and the classical state space models. In essence, state space models compress the context into a smaller state, achieving $O(1)$ time complexity and $O(d)$ space complexity in inference time. State space models are efficient in inference but also limited by the amount of information that can be compressed, i.e. the hidden state size. To mitigate this drawback, [GD23, ] propose a novel selective state space model named Mamba. Mamba  addresses the computation problem of SSM with a selective scan method and corresponding hardware-aware optimization techniques, achieving performance close to transformer-based models of similar sizes while also demonstrating efficiency in training and inference.
@@
RQ: Can Mamba achieve comparable performance to transformer-based language models in the document ranking task?
To comprehensively assess the performance of Mamba, we conduct a rigorous benchmarking study against transformer-based models characterized by varying structures, sizes, pre-training objectives, and attention patterns. Correspondingly, we train neural reranker models following established training methodologies outlined in prior literature [Boy+22, GDC21, Ma+23] This comparison allows us to evaluate the capabilities of different LMs in terms of performance and training efficiency. We find:
@@ 2.1 Structured State Space Sequence Models (S4)
Prior works on state space models [Gu+21, GGB22, Gu+22, SWL22] are linear time invariance (LTI). We use Structured State Space Sequence Models (S4) [GGR21] as a representative example to illustrate the idea behind vanilla state space models.
@@
$h^{\prime}(t)$ $=\textbf{\emph{A}}h(t)+\textbf{\emph{B}}x(t)$ $y(t)$ $=\textbf{\emph{C}}h(t)$ (1)
@@
[…]
Readers might have noticed that Equation 2 looks like RNN while Equation 4 looks like CNN, where $\overline{\textbf{\emph{K}}}$ is a large convolution kernel over the whole input sequence $x$. The parameterization of $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ is independent of input sequence $x$ and is fixed during all time steps, a property referred to as linear time invariance (LTI). Structured state space models (S4) impose structure on the A matrix to be computed efficiently. [^2] An important note is that A matrix also requires a specific initialization method: HiPPO [Gu+20] to memorize the context, the details of which will not be covered in this work. Existing works [GGR21, GGB22, SWL22] mostly deploy diagonal matrix, thus $\textbf{\emph{A}}\in\mathbb{R}^{N\times N}$, $\textbf{\emph{B}}\in\mathbb{R}^{N\times 1}$ and $\textbf{\emph{C}}\in\mathbb{R}^{1\times N}$ matrices are all represented by $N$ parameters.
[…]
@@ 2.2 Selective Scan Structured State Space Sequence Models (S6)
State space models compress context into a smaller hidden state $h_{t}\in\mathbb{R}^{N}$, which limits the model’s effectiveness. [GD23, ] propose to modify a subset of parameters $(\Delta,\textbf{\emph{B}},\textbf{\emph{C}})$ as input-dependent. This modification changes the model from time-invariant to time-varying, and poses challenges to the model’s computational efficiency, as the model now cannot be trained in CNN mode. Specifically, denote the context length as $L$ and batch size as $B$, training in RNN mode yields a total FLOPs of $O(BLDN)$, which cannot be trivially parallelized as each step depends on the output of the previous step.
To tackle the computation problem, [GD23, ] propose a hardware-aware optimization algorithm named Selective Scan. It mainly comprises the following critical improvements:
  • Instead of computing $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}})$ of size $(B,L,D,N)$ in GPU HBM (High-Bandwidth Memory), now SSM parameters $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ are loaded in GPU fast SRAM (Static Random Access Memory) to speed up the computing, and output of size $(B,L,D)$ are written back to HBM. This optimization technique is similar to Flash Attention [Dao24].
  • To avoid the sequential recurrence and to perform the computation in parallel, a work-efficient parallel scan algorithm [Ble, MC18] is used to amortize the computation to multiple threads within GPU.
@@
[GD23, ] parameterize the corresponding selective state space model into a Mamba  block, in which the selective scan structured state space sequence (S6) model only serves as a replacement for Attention component, while the majority of computation is handled by FFN layers and corresponding activation layers. Combining the aforementioned optimization techniques, the parameterized Mamba  block can be computed efficiently and achieve 5× higher throughput on long sequence lengths compared to Transformers of similar sizes, as reported in [GD23].
@@ 3 Methodology and Experiment Setup
Previously we have briefly discussed state space models and the design choices of Mamba  model. In this section, we focus on evaluating Mamba’s efficacy in the document ranking task. To this end, we aim to address the following research questions:
@@ 3.1 Task Formulation
Encoder-decoder models, i.e. T5  family models can be used for feature extraction such as encoder-only models, and can be used for text generation such as decoder-only models. As the encoder part is pre-trained with bidirectional attention, we use the same input format as encoder-only models. Then the FFN layer takes the representation of the first decoded token and predicts the relevance score. This setup is the same as prior works on adapting T5  for document ranking task [Zhu+23, Ni+22].
@@
Table 1: We list the compared backbone language models in terms of size, information, pre-training objective as well as model structure. NSP refers to the next sentence prediction objective.
Model
Size
Information
Pre-training
Objective
Model
Structure
Positional
Encoding
@@
bert-base-uncased
110M
Bidirectional
MLM & NSP
Attention
Learned
bert-large-uncased
330M
Bidirectional
MLM & NSP
Attention
Learned
roberta-base
125M
Bidirectional
MLM
Attention
Learned
roberta-large
355M
Bidirectional
MLM
Attention
Learned
@@
facebook/opt-125m
125M
Unidirectional
CLM
Attention
Learned
facebook/opt-350m
350M
Unidirectional
CLM
Attention
Learned
EleutherAI/pythia-160m
120M
Unidirectional
CLM
Attention
Rotary
EleutherAI/pythia-410m
340M
Unidirectional
CLM
Attention
Rotary
EleutherAI/pythia-1b
860M
Unidirectional
CLM
Attention
Rotary
state-spaces/mamba-130m
130M
Unidirectional
CLM
Mamba
None
state-spaces/mamba-370m
370M
Unidirectional
CLM
Mamba
None
state-spaces/mamba-790m
790M
Unidirectional
CLM
Mamba
None
@@
T5-base
220M
Bidirectional
[…]
Attention
Relative
T5-large
770M
Bidirectional
Denoising
Attention
Relative
  • $\mathbf{a}$ Note that T5  family models are additionally multitask fine-tuned with GLUE tasks [Wan+18].
@@ 3.3 Backbone Language Models
We benchmark the performance of different models on three sizes: models with $\approx$ 110M parameters (e.g. bert-base-uncased); models with $\approx$ 330M parameters (e.g. bert-large-uncased) and models with $>$ 700M parameters (e.g. T5-large and state-spaces/mamba-790m).
For encoder-only models, we choose BERT [Dev+19] and RoBERTa [Liu+19]; for decoder-only models, we include Pythia [Bid+23], OPT [Zha+22] and Mamba [GD23], and we opt for T5 [Raf+20] as the representative encoder-decoder models. The compared models are from different model families, varying in terms of pre-training objectives (e.g. causal language modeling, masked language modeling, denoising, next sentence prediction), information direction (bidirectional vs unidirectional), model structure (Attention vs Mamba) and Positional Encoding (no positional encoding, learned positional encoding, relative positional encoding and rotary positional encoding [Su+24]). An important note is that we do not include instruction-finetuned language models such as Flan-T5 models [Chu+22] for a fair comparison to other language model families. For a review of these models, refer to Table 1.
@@ 3.4 Datasets
Following prior practices [GDC21, Boy+22, Ma+23], for each query and relevant document from the training set, we sample negative documents from the top-100 results returned by a sparse BM25 retriever implemented by Pyserini [^3] https://github.com/castorini/pyserini [Lin+21]. In practice, we fix the random seed and sample 7$\times$2 negative documents for each qrel in the official training set, leading to in total 734,008 training samples, each consisting of 1 relevant document and 7 hard negative documents. Notably, we use the setup recommended by Gao et al. [GDC21] to sample hard negatives from the top-100 results returned by the first-stage retriever (BM25 in our case). The sample training set is used throughout all experiments to ensure fair comparison and reproducibility.
We use the official document ranking Dev set for evaluation. We also use TREC DL19 [Cra+20] and DL20 [Cra+21]. The official Dev set consists of 5,193 queries, with 1 relevant judgment each, while DL19 and DL20 consist of 43 and 45 queries respectively with denser relevance judgments.
We use the reranking setup in evaluation as prior practices [Boy+22, GDC21, Ma+23] where a first-stage retriever BM25 retrieves top-100 documents and our trained reranker models perform the reranking. We report Mean Reciprocal Rank@100 for Dev set and NDCG@10 for DL19 and DL20 to be consistent with the official evaluation. Deploying better first-stage retrievers can lead to better reranking performance but is beyond the scope of this study.
@@ 3.5 Training Details and Hyperparameters
We truncate the input to 512 tokens to fit in the context window of BERT  and RoBERTa and T5  models (also referred to as FirstP by [DC19, ]). For autoregressive LMs (OPT , Pythia  and Mamba), although they have a longer context length, we also use 512 tokens to keep a fair comparison. We use the official implementation of Mamba  models [^4] https://github.com/state-spaces/mamba and weights of other models are downloaded from Huggingface. [^5] https://huggingface.co/models
For all experiments, we optimize the reranker model with AdamW optimizer [LH19], with betas=(0.9, 0.999), eps=1e-08, weight_decay=0.01. In our preliminary experiments, we notice models larger than 300M can suffer from overfitting, therefore we use a smaller learning rate for larger models. We use a linear warmup scheduling where we gradually increase the learning rate in the first 1000 training steps and decrease to zero until the end of training. We use PyTorch native mixed precision training and Flash Attention 2 [Dao24] for higher throughput. For models larger than 700M, i.e. T5-large, EleutherAI/pythia-1b and state-spaces/mamba-790m, we use LoRA [Hu+21] to reduce memory consumption and speed up training.
We train the reranker models for 1 single epoch, as recommended by Ma et al. [Ma+23]. Through the course of training, for each reranker model, we save 10 intermediate checkpoints to monitor the training trajectory and report the performance of the final checkpoint. Refer to Table 2 for a detailed listing of hyperparameters and training details.

### v3

@@ Abstract
Transformer structure has achieved great success in multiple applied machine learning communities, such as natural language processing (NLP), computer vision (CV) and information retrieval (IR). Transformer architecture’s core mechanism — attention requires $O(n^{2})$ time complexity in training and $O(n)$ time complexity in inference. Many works have been proposed to improve the attention mechanism’s scalability, such as Flash Attention and Multi-query Attention. A different line of work aims to design new mechanisms to replace attention. Recently, a notable model structure — Mamba, which is based on state space models, has achieved transformer-equivalent performance in multiple sequence modeling tasks. In this work, we examine Mamba’s efficacy through the lens of a classical IR task — document ranking. A reranker model takes a query and a document as input, and predicts a scalar relevance score. This task demands the language model’s ability to comprehend lengthy contextual inputs and to capture the interaction between query and document tokens. We find that (1) Mamba models achieve competitive performance compared to transformer-based models with the same training recipe; (2) but also have a lower training throughput in comparison to efficient transformer implementations such as flash attention. We hope this study can serve as a starting point to explore Mamba models in other classical IR tasks. Our code implementation is made public to facilitate reproducibility. Refer to Xu et al. (2025d) for more comprehensive experiments and results, including passage ranking.
@@ 1 Introduction
Transformer structure has been the status quo within NLP and IR community since the introduction of Transformers (Vaswani et al., 2017) and BERT (Devlin et al., 2019). Compared to the classical RNN structure, the transformer structure is better at capturing long-range dependencies and also can be pre-trained on a large scale. However, in inference time, the transformer structure requires $O(n)$ time complexity and $O(n\cdot d)$ space complexity to store KV-cache, where $n$ denotes sequence length and $d$ denotes the size of the hidden state, making it less efficient compared to RNN. In addition, the transformer structure is also hard to extrapolate to context length longer than its pre-training length.
Recently, there has been a growing interest in developing alternative structures for modeling sequence data. For example, RWKV (Peng et al., 2023) combines the efficient parallelizable training of transformers with the efficient inference of RNNs. Another notable structure is state space models (SSM) (Gu and Dao, 2023), which are related to CNN, RNN and the classical state space models. In essence, state space models compress the context into a smaller state, achieving $O(1)$ time complexity and $O(d)$ space complexity in inference time. State space models are efficient in inference but also limited by the amount of information that can be compressed, i.e. the hidden state size. To mitigate this drawback, Gu and Dao (2023) propose a novel selective state space model named Mamba. Mamba   addresses the computation problem of SSM with a selective scan method and corresponding hardware-aware optimization techniques, achieving performance close to transformer-based models of similar sizes while also demonstrating efficiency in training and inference.
@@
RQ: Can state space models (e.g., Mamba) achieve comparable performance to transformer-based language models in the document ranking task?
To comprehensively assess the performance of Mamba, we conduct a rigorous benchmarking study against transformer-based models characterized by varying structures, sizes, pre-training objectives, and attention patterns. Correspondingly, we train neural reranker models following established training methodologies outlined in prior literature (Boytsov et al., 2022; Gao et al., 2021; Ma et al., 2023) This comparison allows us to evaluate the capabilities of different LMs in terms of performance and training efficiency. We find:
@@ 2.1 Structured State Space Sequence Models (S4)
Prior works on state space models (Gu et al., 2021b; Gupta et al., 2022; Gu et al., 2022; Smith et al., 2022) are linear time invariance (LTI). We use Structured State Space Sequence Models (S4) Gu et al. (2021a) as a representative example to illustrate the idea behind vanilla state space models.
@@
[…]
@@
$\overline{\textbf{\emph{K}}}$ $=(\textbf{\emph{C}}\overline{\textbf{\emph{B}}},C\overline{\textbf{\emph{A}}\textbf{\emph{B}}},\ldots,C\overline{\textbf{\emph{A}}}^{n-1}\overline{\textbf{\emph{B}}})$ $y$ $=x*\overline{\textbf{\emph{K}}}$ (4)
Readers might have noticed that Equation 2 looks like RNN while Equation 4 looks like CNN, where $\overline{\textbf{\emph{K}}}$ is a large convolution kernel over the whole input sequence $x$. The parameterization of $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ is independent of input sequence $x$ and is fixed during all time steps, a property referred to as linear time invariance (LTI). Structured state space models (S4) impose structure on the A matrix to be computed efficiently. [^1] An important note is that A matrix also requires a specific initialization method: HiPPO Gu et al. (2020) to memorize the context, the details of which will not be covered in this work. Existing works (Gu et al., 2021a; Gupta et al., 2022; Smith et al., 2022) mostly deploy diagonal matrix, thus $\textbf{\emph{A}}\in\mathbb{R}^{N\times N}$, $\textbf{\emph{B}}\in\mathbb{R}^{N\times 1}$ and $\textbf{\emph{C}}\in\mathbb{R}^{1\times N}$ matrices are all represented by $N$ parameters.
Equation 1-4 can be generalized to $D$-channel features, i.e. $x_{t},y_{t}\in\mathbb{R}^{D}$, a concrete example might be a word embedding of dimension $D$. In this case, computation of $\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}}$ is applied to each channel independently.
@@ 2.2 Selective Scan Structured State Space Sequence Models (S6)
State space models compress context into a smaller hidden state $h_{t}\in\mathbb{R}^{N}$, which limits the model’s effectiveness. Gu and Dao (2023) propose to modify a subset of parameters $(\Delta,\textbf{\emph{B}},\textbf{\emph{C}})$ as input-dependent. This modification changes the model from time-invariant to time-varying, and poses challenges to the model’s computational efficiency, as the model now cannot be trained in CNN mode. Specifically, denote the context length as $L$ and batch size as $B$, training in RNN mode yields a total FLOPs of $O(BLDN)$, which cannot be trivially parallelized as each step depends on the output of the previous step.
To tackle the computation problem, Gu and Dao (2023) propose a hardware-aware optimization algorithm named Selective Scan. It mainly comprises the following critical improvements:
  • Instead of computing $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}})$ of size $(B,L,D,N)$ in GPU HBM (High-Bandwidth Memory), now SSM parameters $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ are loaded in GPU fast SRAM (Static Random Access Memory) to speed up the computing, and output of size $(B,L,D)$ are written back to HBM. This optimization technique is similar to Flash Attention Dao (2024).
  • To avoid the sequential recurrence and to perform the computation in parallel, a work-efficient parallel scan algorithm Blelloch ; Martin and Cundy (2018) is used to amortize the computation to multiple threads within GPU.
@@
Gu and Dao (2023) parameterize the corresponding selective state space model into a Mamba block, in which the selective scan structured state space sequence (S6) model only serves as a replacement for Attention component, while the majority of computation is handled by FFN layers and corresponding activation layers. Combining the aforementioned optimization techniques, the parameterized Mamba block can be computed efficiently and achieve 5× higher throughput on long sequence lengths compared to Transformers of similar sizes, as reported in Gu and Dao (2023).
@@ 3 Methodology and Experiment Setup
We have briefly discussed state space models and the design choices of Mamba model. In this section, we focus on evaluating Mamba’s efficacy in the document ranking task. To this end, we aim to address the following research questions:
@@ 3.1 Task Formulation
Encoder-decoder models, i.e. T5   family models can be used for feature extraction such as encoder-only models, and can be used for text generation such as decoder-only models. As the encoder part is pre-trained with bidirectional attention, we use the same input format as encoder-only models. Then the FFN layer takes the representation of the first decoded token and predicts the relevance score. This setup is the same as prior works on adapting T5 for the document ranking task (Zhuang et al., 2023; Ni et al., 2022).
@@
Table 1: We list the compared backbone language models in terms of size, information, pre-training objective as well as model structure. NSP refers to the next sentence prediction objective. $\mathbf{a}$ Note that T5   family models are additionally multitask fine-tuned with GLUE tasks Wang et al. (2018).
Model Size Information Pre-training Objective Model Structure Positional Encoding
@@
bert-base-uncased 110M Bidirectional MLM & NSP Attention Learned
bert-large-uncased 330M Bidirectional MLM & NSP Attention Learned
roberta-base 125M Bidirectional MLM Attention Learned
roberta-large 355M Bidirectional MLM Attention Learned
@@
facebook/opt-125m 125M Unidirectional CLM Attention Learned
facebook/opt-350m 350M Unidirectional CLM Attention Learned
EleutherAI/pythia-160m 120M Unidirectional CLM Attention Rotary
EleutherAI/pythia-410m 340M Unidirectional CLM Attention Rotary
EleutherAI/pythia-1b 860M Unidirectional CLM Attention Rotary
state-spaces/mamba-130m 130M Unidirectional CLM Mamba None
state-spaces/mamba-370m 370M Unidirectional CLM Mamba None
state-spaces/mamba-790m 790M Unidirectional CLM Mamba None
@@
T5-base 220M Bidirectional Denoisinga Attention Relative
T5-large 770M Bidirectional Denoising Attention Relative
@@ 3.3 Backbone Language Models
We benchmark the performance of different models on three sizes: models with 110M parameters (e.g. bert-base-uncased); models with 330M parameters (e.g. bert-large-uncased) and models with >> 700M parameters (e.g. T5-large and state-spaces/mamba-790m).
For encoder-only models, we choose BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019); for decoder-only models, we include Pythia (Biderman et al., 2023), OPT (Zhang et al., 2022) and Mamba (Gu and Dao, 2023), and we opt for T5 (Raffel et al., 2020) as the representative encoder-decoder models. The compared models are from different model families, varying in terms of pre-training objectives (e.g. causal language modeling, masked language modeling, denoising, next sentence prediction), information direction (bidirectional vs unidirectional), model structure (Attention vs Mamba) and Positional Encoding (no positional encoding, learned positional encoding, relative positional encoding and rotary positional encoding (Su et al., 2024)). An important note is that we do not include instruction-finetuned language models such as Flan-T5 models (Chung et al., 2022) for a fair comparison to other language model families. For a review of these models, refer to Table 1.
@@ 3.4 Datasets
Following prior practices (Gao et al., 2021; Boytsov et al., 2022; Ma et al., 2023), for each query and relevant document from the training set, we sample negative documents from the top-100 results returned by a sparse BM25 retriever implemented by Pyserini [^2] https://github.com/castorini/pyserini (Lin et al., 2021). In practice, we fix the random seed and sample 72 negative documents for each qrel in the official training set, leading to in total 734,008 training samples, each consisting of 1 relevant document and 7 hard negative documents. Notably, we use the setup recommended by Gao et al. (2021) to sample hard negatives from the top-100 results returned by the first-stage retriever (BM25 in our case). The sample training set is used throughout all experiments to ensure fair comparison and reproducibility.
We use the official document ranking Dev set for evaluation. We also use TREC DL19 (Craswell et al., 2020) and DL20 (Craswell et al., 2021). The official Dev set consists of 5,193 queries, with 1 relevant judgment each, while DL19 and DL20 consist of 43 and 45 queries respectively with denser relevance judgments.
We use the reranking setup in evaluation as prior practices (Boytsov et al., 2022; Gao et al., 2021; Ma et al., 2023) where a first-stage retriever BM25 retrieves top-100 documents and our trained reranker models perform the reranking. We report Mean Reciprocal Rank@100 for Dev set and NDCG@10 for DL19 and DL20 to be consistent with the official evaluation. Deploying better first-stage retrievers can lead to better reranking performance but is beyond the scope of this study.
@@ 3.5 Training Details and Hyperparameters
We truncate the input to 512 tokens to fit in the context window of BERT and RoBERTa and T5 models (also referred to as FirstP by Dai and Callan (2019)). For autoregressive LMs (OPT, Pythia and Mamba), although they have a longer context length, we also use 512 tokens to keep a fair comparison. We use the official implementation of Mamba models [^3] https://github.com/state-spaces/mamba and weights of other models are downloaded from Huggingface. [^4] https://huggingface.co/models
For all experiments, we optimize the reranker model with AdamW optimizer (Loshchilov and Hutter, 2019), with betas=(0.9, 0.999), eps=1e-08, weight_decay=0.01. In our preliminary experiments, we noticed models larger than 300M can suffer from overfitting, therefore we use a smaller learning rate for larger models. We use a linear warmup scheduling where we gradually increase the learning rate in the first 1000 training steps and decrease to zero until the end of training. We use PyTorch native mixed precision training and Flash Attention 2 Dao (2024) for higher throughput. For models larger than 700M, i.e. T5-large, EleutherAI/pythia-1b and state-spaces/mamba-790m, we use LoRA (Hu et al., 2021) to reduce memory consumption and accelerate training.
We train the reranker models for 1 single epoch, as recommended by Ma et al. (2023). Through the course of training, for each reranker model, we save 10 intermediate checkpoints to monitor the training trajectory and report the performance of the final checkpoint. Refer to Table 2 for a detailed listing of hyperparameters and training details.